### v2

# Trapped-ion Laser Cooling in Structured Light Fields

Abstract
Laser cooling is fundamental to quantum computation and metrology with trapped ions, and can occupy a majority of runtime in current systems. A key limitation to cooling arises from unwanted carrier excitation, which in typically used running wave (RW) fields invariably accompanies the sideband transitions effecting cooling. We consider laser cooling in structured light profiles enabling selective sideband excitation with nulled carrier drive; motivated by integrated photonic approaches’ passive phase and amplitude stability, we propose simple configurations realizable with waveguide addressing using either standing wave (SW) or first-order Hermite-Gauss (HG) modes. We quantify performance of Doppler cooling from beyond the Lamb-Dicke regime (LDR), and ground-state (GS) cooling using electromagnetically induced transparency (EIT) leveraging these field profiles. Carrier-free EIT offers significant benefits simultaneously in cooling rate, motional frequency bandwidth, and final phonon number. Carrier-free Doppler cooling’s advantage is significantly compromised beyond the LDR but continues to hold, indicating such configurations are applicable for highly excited ions. Our simulations focus on level structure relevant to ${}^{40}\text{Ca}^{+}{}$, though the carrier-free configurations can be generally applied to other species. We also quantify performance limitations due to polarization and modal impurities relevant to experimental implementation. Our results indicate potential for simple structured light profiles to alleviate bottlenecks in laser cooling, and for scalable photonic devices to improve basic operation quality in trapped-ion systems.
preprint: APS/123-QED
Trapped ions constitute a leading platform for quantum computation (QC) Bruzewicz et al. (2019) and simulation, with recent systems demonstrating algorithms on systems of tens of qubits with arbitrary connectivity Moses et al. (2023); Postler et al. (2024), leveraging the high-fidelity entangling gates Ballance et al. (2016); Gaebler et al. (2016); Clark et al. (2021); Löschnauer et al. (2024), state preparation, measurement, and readout Harty et al. (2014) achievable in this platform. The high coherence and precise control of systematics achievable in ion traps also enables timekeeping at the present limits of accuracy Brewer et al. (2019); Hausser et al. (2024).
High-fidelity trapped-ion control in QC systems generally requires laser cooling to near the motional ground state (GS) Leibfried et al. (2003), also of interest for next-generation clocks Kulosa et al. (2023). To this end, Doppler cooling to near the Doppler limit may be followed by electromagnetically induced transparency (EIT) cooling Morigi et al. (2000a); Roos et al. (2000); Eschner et al. (2003), polarization gradient cooling (PGC) Birkl et al. (1994), and/or sideband cooling (SBC) Diedrich et al. (1989), to successively bring expectation phonon number $\bar{n}$ to near the GS Leibfried et al. (2003). Due to inevitable heating of atoms during operation De Leon et al. (2021) and because readout and transport operations repeated in trapped-ion QC operation Kielpinski et al. (2002) can excite ion motion, laser cooling generally must be repeatedly applied. Given the cooling rates achievable, laser cooling can occupy a majority of runtime in current trapped-ion QC systems Pino et al. (2021). Since operation times are a key limitation of trapped-ion systems, cooling schemes with improved cooling rate and final phonon number may address a crucial bottleneck in trapped-ion systems.
Figure 1: (a) Steady state excitation population of a two-level system as a function of detuning $\Delta$. Carrier, BSB, and RSB transitions correspond to black, blue, and red dotted lines respectively. Due to power broadening, higher beam power ($s_{0}=2\Omega_{0}^{2}/\Gamma^{2}$) reduces the contrast between the RSB and BSB scattering rates, compromising cooling performance. (b) Probability distribution of a ${}^{40}\text{Ca}^{+}$ ion in thermal states for $\omega_{m}=2\pi\times 3$ MHz, centered at the nodal point of a $\lambda=397$ nm SW and a $w_{0}=500$ nm TE10 mode. The wave function of higher phonon number samples a larger field region, suggesting different cooling behaviors when outside/inside LDR. (c) Two experimental schemes for carrier-suppressed driving using integrated photonics, based on two counter-propagating beams from the same input to form a passively phase-stable SW, or a TE10 HG mode delivered from a single grating emitter fed by a TE1 guided mode Beck et al. (2024). The cooling beam’s intensity profiles are sketched on the right. (d) The eight-level structure relevant to ${}^{40}\text{Ca}^{+}$ ($S_{1/2},P_{1/2},D_{3/2}$) with lasers for Doppler cooling. $\Delta_{c},\Delta_{r}$ are detuning between $S_{1/2}$ and $P_{1/2}$, $D_{3/2}$ and $P_{1/2}$ respectively at zero magnetic field.
Key limitations to cooling performance arise from couplings that accompany the desired interactions when employing the usual plane-wave-like running wave (RW) fields. In a frame rotating with laser frequency $\omega$, interaction between a RW laser field and an atom represented as a two-level system with ground and excited states $|g\rangle$ and $|e\rangle$, oscillating in a single motional mode at angular frequency $\omega_{m}$ can be described with a dipole interaction Hamiltonian expanded as:
$\langle e|\hat{V}_{\mathrm{dip}}|g\rangle=\frac{\Omega_{0}}{2}e^{i\mathbf{k}\cdot\mathbf{\hat{R}}}=\frac{\Omega_{0}}{2}(1+i\eta(\hat{a}+\hat{a}^{\dagger})+...),$ (1)
for wave vector $\mathbf{k}$ with $\left\lvert\mathbf{k}\right\rvert=k_{0}$ in free space, Rabi frequency $\Omega_{0}$, position operator $\mathbf{\hat{R}}=\mathbf{x_{0}}(\hat{a}+\hat{a}^{\dagger})$, and LD parameter $\eta=\eta_{0}\cos(\theta)$ where $\eta_{0}=k_{0}x_{0}=k_{0}\sqrt{\frac{\hbar}{2m\omega_{m}}}$ and $\theta$ is the angle between $\mathbf{k}$ and $\mathbf{x_{0}}$. The second term $\propto\eta$ describes coupling to motional sidebands due to first-order field gradients desired for cooling; larger scattering rate on the red sideband (RSB) as compared to the blue sideband (BSB) leads to cooling (Fig. 1). The first order term, due to the field at the ion location, represents carrier coupling that does not affect the motional state. Since typically $\eta<1$, the carrier drive strength exceeds the sideband and is problematic for two reasons. As $\Omega_{0}$ is increased to cool more rapidly, this carrier excitation saturates the internal population, reducing contrast in the excitation rate between the RSB and BSB, thereby limiting cooling rates and compromising the final phonon number achievable (Fig. 1a). Additionally, while carrier excitation is associated with no motional excitation, the subsequent spontaneous decay results in net heating. Thus, carrier excitation limits achievable cooling rate and additionally causes net heating.
These limitations can be alleviated via suppression of carrier excitation and selective sideband coupling achievable at appropriate locations in structured light fields Cirac et al. (1992); Zhang et al. (2012); deLaubenfels et al. (2015); Vasquez et al. (2023). For electric dipole transitions within atoms trapped at points with zero intensity and nonzero field gradient, e.g. at SW nodes or nodal lines of higher-order Hermite-Gaussian (HG) modes (Fig. 1b), the carrier term (and all even orders of the expansion in Eq. 1) is nulled, with the first sideband term the leading contribution.
Recent development of integrated waveguide photonics for optical control in planar ion traps Mehta et al. (2016); Niffenegger et al. (2020); Mehta et al. (2020); Vasquez et al. (2023); Ivory et al. (2021); Kwon et al. (2024); Mordini et al. (2024) offers a route to realize these mechanisms without the phase drifts and pointing instability that pose significant challenges in conventional free-space implementations Schmiegelow et al. (2016). For example, SW generation straightforwardly leverages counterpropagating beams derived from a single waveguide input (Fig. 1c, top) as demonstrated in Vasquez et al. (2023), and first-order HG mode generation can be achieved via emission of higher-order waveguide modes, e.g. a TE10 mode Beck et al. (2024) allowing use of transverse field gradients at the intensity node along the beam center (Fig. 1c, bottom). While the form of the interaction Hamiltonian is the same to lowest order for these two cases, the effective LD parameter in the latter case, $\eta_{\text{HG}}=\frac{2\sqrt{2}}{w_{0}}x_{0}$ West et al. (2021) reflects motional coupling from transverse field gradients, additionally offering avenues to engineer interactions to accommodate different ion crystal configurations and motional mode orientations. While we focus on these relatively simple field configurations, similar considerations would apply to more general structured light profiles Verde et al. (2023).
In this work, we examine the practical performance of Doppler and EIT cooling using structured light fields through master equation simulation. We consider ${}^{40}\text{Ca}^{+}$’s relevant eight-level structure ($S_{1/2},P_{1/2},D_{3/2}$) in Fig. 1d, although the methods analyzed are applicable to other species. Our treatment extends previous theoretical work in considering Doppler cooling from well beyond the LDR (Fig. 1b), and accounting for sideband saturation effects.
In the results presented below we show significant enhancement possible for carrier-free EIT GS cooling, simultaneously in cooling rate, motional mode frequency bandwidth, and over an order of magnitude reduction in final phonon number $\bar{n}_{ss}$ achievable. We quantify the sensitivity to these enhancements to polarization impurities, a likely experimental limitation. Realization of these simultaneous advantages may significantly reduce the SBC required to achieve GS occupancies required for high-fidelity control. For Doppler cooling, despite the expected degradation in SW cooling of highly excited ions as compared to within the LDR, we still find that a modest enhancement for SW cooling obtains. While sub-Doppler cooling via PGC also can leverage integrated photonic beam delivery Hattori et al. (2022) and benefit from ion positioning at polarization SW nodes, the expected enhancement in $\bar{n}_{ss}$ compared to with RWs is 2×2$\times$2 × Joshi et al. (2020). Given the more significant gains predicted for EIT here, we focus on EIT as a sub-Doppler and GS cooling method. We present our results for the simpler Doppler followed by EIT cooling below, commenting on the impact of $\eta_{0}$ on both schemes. This work indicates potential for scalable photonic hardware to alleviate bottlenecks in trapped-ion control, focusing here on laser cooling, and informs experimental implementation in integrated devices.
I Carrier-free Doppler cooling
The treatment of Cirac et al. (1992) showed that Doppler cooling for ions within the LDR at SW nodes allows higher cooling rates via use of higher cooling beam intensities, as well as lower final $\bar{n}_{ss}$, than possible with RWs. Since SW cooling is limited by saturation on sideband transitions ($\Omega_{\mathrm{SB}}\propto\eta\Omega_{0}$) instead of the carrier ($\Omega_{\mathrm{car}}=\Omega_{0}$) as in the usual RW, in a SW configuration and for ions within the LDR, $W_{c}\propto\Omega_{\mathrm{SB}}^{2}$ can be increased by of order ${\sim 1}/\eta^{2}$ before saturating the two-level system. This delayed saturation enables a significant gain in $W_{c}$ given that typically, $\eta\sim 0.1$. Furthermore, an approximately 2×2$\times$2 × reduction in final phonon number $\bar{n}_{ss}$ is predicted.
However, significantly different behavior, including heating, appears at other positions in the SW profile Cirac et al. (1992). Wavefunctions of highly excited ions outside the LDR, of practical interest to Doppler cooling, span lengthscales comparable to the wavelength (Fig. 1b), and hence compromised SW cooling performance is expected as compared to within the LDR as treated in Cirac et al. (1992). To study the cooling behavior beyond the LDR, here we employ numerical Lindblad master equation simulations including higher-order expansion terms (see Appendix). Our simulation quantifies the enhancement of Doppler cooling using SWs as compared to RWs both outside and inside the LDR, including effects of saturation from sideband excitation. We also discuss experimental configurations of carrier-free Doppler cooling based on integrated photonic addressing Vasquez et al. (2023); Mehta et al. (2020).
In particular, we simulate the eight-level ${}^{40}\text{Ca}^{+}$ structure as an example with a $\hat{\pi}$-polarized cooling beam ($\Omega_{c}$, $\Delta_{c}$) addressing the $S_{1/2}\leftrightarrow P_{1/2}$ transition and a repumper ($\Omega_{r}=2\pi\times 10$ MHz, $\Delta_{r}$) with polarization $\hat{\epsilon}_{p}=\hat{\pi}/\sqrt{2}+(\hat{\sigma}_{+}-\hat{\sigma}_{-})/2$ between $D_{3/2}\leftrightarrow P_{1/2}$ as shown in Fig. 1a, and with a quantizing magnetic field of $B=1$ G. Time evolution of the expectation phonon number $\bar{n}(t)$ starting from a thermal state $\bar{n}(0)$ is extracted from master equation evolution. The cooling rate at a given $\bar{n}$ is defined as $W_{c}=-\dot{\bar{n}}/\bar{n}$, and the cooling limit $\bar{n}_{ss}$ the final steady-state expectation phonon number.
Figure 2: (a) Cooling rate (line) and limit (point) for $\omega_{m}=2\pi\times 3$ MHz and $\Delta_{c}=-\Gamma/2$. EF$=1.5$ for $n=50$ (solid line), EF$=4.9$ for $n=5$ (dotted line). The additional D states decrease the cooling for both SW and RW but also reduce the RW carrier heating. (b) Time evolution of SW and RW cooling for $\omega_{m}=2\pi\times 3$ MHz. Using $\Delta_{c}=-0.5\Gamma$ and starting at $\bar{n}=50$, SW Doppler cooling reaches $\bar{n}=4$ within $34~{}\mu s$ with optimal Rabi frequency $\Omega_{sw}=1.8\Gamma$, while RW Doppler cooling takes $75~{}\mu s$ with optimal $\Omega_{rw}=0.8\Gamma$.
Fig. 2a shows $W_{c}$ in the full eight-level system as a function of $\Omega_{c}$ for $\omega_{m}=2\pi\times 3$ MHz. $\Delta_{c}$ is set to $-\Gamma_{c}/2$ which optimizes the cooling limit $\bar{n}_{ss}$. For small $\Omega_{0}$, RW and SW Doppler cooling rates increase quadratically with $\Omega_{0}$ both inside ($\bar{n}=5$) and outside the LDR ($\bar{n}=50$) before saturation. For higher $\bar{n}$, the EF drops but remains larger than 1, achieved around the same $\Omega_{c}$ both outside and near the LDR. Combined with the consistent improvement in $\bar{n}_{ss}$, this validates the advantage of SW Doppler cooling in the experiments. In Fig. 2b, we simulate a complete Doppler cooling process from $\bar{n}=50$ outside the LDR to the Doppler limit, using the optimal parameters for cooling rate found in Fig. 2a. We observe SW Doppler cooling take less than half of the time as compared to RW to reach $\bar{n}=4$, a reasonable starting point for ground-state cooling.
Experimental implementation of carrier-free Doppler cooling may employ the configurations shown in Fig. 1c. Though for practical beam waists, the TE10’s effective $\eta$ is lower than in a RW, optimal sideband coupling can be recovered by increasing $\Omega_{0}$ inside the LDR to hold $\eta\Omega_{0}$ constant. This allows cooling radial modes of multi-ion strings without requiring precise positioning of multiple ions along a series of SW fringes. Outside the LDR, however, cooling through higher-order sideband excitations (with coupling strengths proportional to $\eta^{k}$ for a $k$th order sideband) contribute significantly to cooling, which are attenuated for lower driving beam LD parameter $\eta$ and fixed first-order coupling $\eta\Omega$. In contrast the heating through spontaneous decay on a $k$th order sideband is proportional to $\eta_{0}^{k}$ which is independent of the driving beam profile. As a result, the EF in cooling rate achievable with SW cooling decreases outside the LDR when using lower $\eta$ values. For example, in our eight-level system simulations we observe that the EF for SW cooling with $\eta=\eta_{0}/4$ compared to the best case RW with $\eta=\eta_{0}$ drops to nearly unity for $\bar{n}=50$.
Our results quantify the performance of carrier-free Doppler cooling beyond the LDR previously treated. Though using ${}^{40}\text{Ca}^{+}$ in the simulation as an example, similar advantage would hold for other ion species and are more favorable for lower intrinsic $\eta_{0}$ Xing and Mehta (2023). While the advantage offered is modest in comparison to well within the LDR, our results indicate that such configurations can still be applied for highly excited ions. Therefore, early cooling stages can share the same cooling beam profile with carrier-free GS cooling schemes, for which more substantial enhancements are predicted.
II Carrier-free GS EIT Cooling
Figure 3: (a) Level structure and fields relevant to EIT cooling of ${}^{40}\text{Ca}^{+}$. $\Delta_{p},\Delta_{c}$ are detunings of pump and cooling laser fields coupling $|S_{1/2},m_{j}=-1/2\rangle\leftrightarrow|P_{1/2},m_{j}=1/2\rangle$ and $|S_{1/2},m_{j}=1/2\rangle\leftrightarrow|P_{1/2},m_{j}=1/2\rangle$ respectively. $\Delta_{r}$ is the repumper’s detuning from $\text{D}_{3/2}$ and $P_{1/2}$ at $B=0$. (b) Excitation spectrum as a function of the cooling beam’s detuning, for fixed $\Delta_{p}$. Insets show choice of $\Delta_{c}$ and AC stark shift $\delta$ relative to the motional frequency $\omega_{m}$ to be cooled, for both RW and SW implementations. (c) Comparison of cooling limit over a range of motional frequencies between simulation results (points) and analytical calculation (solid line) for low-intensity $\Omega_{c}$ beam. $\Omega_{p}$ and $\Delta_{c}$ are optimized for $\omega_{m0}=2\pi\times 3$ MHz and $\Delta_{p}=5\Gamma$. Triangular and circular dots correspond to RW and SW pump beams, respectively. (d) Experimental schematics of carrier-free EIT cooling using integrated photonics. The first builds on the design for Doppler cooling with a similar integrated SW cooling beam and free-space RW pump beam. The second uses integrated photonics for all beams. A vertically emitted pump beam may achieve high circular polarization purity Massai et al. (2022) relative to a vertically oriented $B$-field. The cooling beam emitted in TM10 has a SW-like beam profile and comes from a shallow angle to satisfy the polarization requirements discussed below. Motional modes with normal mode vector projection along the cooling axis labeled would be cooled in either case.
EIT-based cooling generally requires a pump beam (with Rabi frequency $\Omega_{p}$) inducing atomic excitation undergoing quantum interference with a weaker cooling beam ($\Omega_{c}$) in an effective three-level system (Fig. 3a). This results in a Fano profile in the cooling beam’s excitation spectrum, with an absorption null at $\Delta_{c}=\Delta_{p}$ and a nearby bright peak separated by the AC stark shift $\delta\equiv\frac{1}{2}\sqrt{\Omega_{p}^{2}+\Delta_{p}^{2}}-\Delta_{p}$ induced by the pump beam Morigi et al. (2000b); Roos et al. (2000), as shown in Fig. 3b. In the usual configuration with RW optical fields, $\Delta_{c}$ is tuned to the EIT null to suppress carrier excitation, and $\delta$ is matched to the frequency $\omega_{m}$ of the motional mode to be cooled, such that its RSB is preferentially driven (Fig. 3b). The high contrast between the RSB and the BSB excitation compared to that achieved with typical dipole transitions employed in Doppler cooling gives rise to cooling with a low final state number Morigi et al. (2000a); Roos et al. (2000); Lechner et al. (2016). However the nonzero BSB excitation amplitude limits final phonon number achievable, and the Fano profile is usually derived for weak probe beam, and fails to describe actual dynamics for large $\Omega_{c}$.
In a carrier-free configuration, the ion can be positioned at a zero-intensity nodal point of the $\Omega_{c}$ beam. The internal state dressing fundamental to EIT relies on intensity of the pump beam $\Omega_{p}$, which can be delivered either as a RW or as a SW with anti-node at the ion location, with comparable performance. With the carrier transition nulled by the spatial coherence of the SW $\Omega_{c}$ field, carrier-free EIT utilizes the pump beam with adjusted parameters to null the BSB excitation amplitude. As shown in Fig. 3b, the cooling beam’s detuning is chosen such that the BSB and RSB lie at the EIT null and peak, respectively. That is, $\Delta_{c}=\Delta_{p}+\omega_{m0}$ and $\delta=2\omega_{m0}$. As in resolved sideband cooling, carrier-free EIT in the LDR then in principle drives strictly the RSB only, thus allowing extremely low final $\bar{n}_{\text{ss}}$ Zhang et al. (2012). Additionally, due to suppression of carrier coupling by the cooling beam, higher $\Omega_{c}$ can be applied to realize faster cooling as compared to the standard RW EIT. Additionally, carrier-free EIT can be expected to offer a larger bandwidth of motional mode frequencies cooled, due to the larger range of detunings from $\Delta_{c}$ for which high contrast between the RSB and BSB excitation probabilities obtains (see Fig. 3b). Thus as compared to RW-based EIT cooling, carrier-free EIT cooling is expected to enable GS cooling of multiple motional modes, simultaneously to lower final phonon numbers. Fig. 3c shows calculated cooling limit in an approximate three-level system for the same pump beam detuning $\Delta_{p}=5\Gamma$ and center motional frequency $\omega_{m0}=2\pi\times 3$ MHz, as a function of motional frequency $\omega_{m}$. Under the low cooling beam intensity limit, $\bar{n}_{ss}$ is independent of $\Omega_{c}$ and whether the pump beam has RW or SW field profile. indicating both the agreement between analytic treatment (lines, and see Appendix) and master-equation simulation (points), along with the significantly lower $\bar{n}_{ss}$ achievable over broader bandwidths using the carrier-free configuration.
The analytical treatment described in the Appendix accurately describes cooling rate and limit for a three-level model for weak $\Omega_{c}$, as validated against master equation simulation in Fig. 3c. To describe the cooling behavior with large $\Omega_{c}$ near the saturation limit where the analytical treatment breaks down, we focus on master equation simulations of the relevant eight-level system relevant to ${}^{40}\text{Ca}^{+}$ at $B=10$ G and examine the achievable cooling rate, phonon number limit, and bandwidth. Theree different couplings are present in the system: a $\hat{\sigma}_{+}$-polarized pump beam $\Omega_{p}$ between $|S_{1/2},m_{j}=-1/2\rangle$ and $|P_{1/2},m_{j}=+1/2\rangle$, a $\hat{\pi}$-polarized cooling beam $\Omega_{c}$ between $|S_{1/2},m_{j}=+1/2\rangle$ and $|P_{1/2},m_{j}=+1/2\rangle$, and a resonant $\hat{\epsilon}_{p}$-polarized repumper $\Omega_{r}$ between the $D_{3/2}$ and $P_{1/2}$ states. Similar to Doppler cooling inside the LDR, the carrier’s relative strength is stronger for smaller $\eta$, suggesting larger possible advantage for low $\eta$. However, carrier-free EIT’s EF in cooling rate is not strictly proportional to $\eta_{0}^{-2}$ because the carrier transition in the RW EIT has also been nulled to first order. Motivated by potential integrated geometries for beam delivery, we also quantify the dependence of these enhancements on imperfect polarization.
Two practical candidate beam geometries for carrier-nulled EIT cooling are shown in Fig. 3d, based either on SW or HG delivery of the cooling beam. Motivated by the full level structure relevant to ${}^{40}\text{Ca}^{+}{}$ (Fig. 3a), we consider geometries to best satisfy the polarization requirement of $\Omega_{c}$ and $\Omega_{p}$. As detailed in the Appendix, in either case, the motional modes cooled are those with some projection along the cooling beam’s field gradient; in contrast to RW-based cooling, the gradient of the $\Omega_{p}$ beam does not enter to leading order, and hence both the SW or HG direction can be tuned with respect to the trap axis to cool a combination of axial and radial modes.
Figure 4: 8L EIT cooling scan of RW (left) and SW (right) on $\Omega_{c}$ and $\omega_{m}$. The first and second rows showcase the cooling rate and final state phonon number respectively. $\Omega_{p}=2\pi\times 52.2(36.4)$ MHz and $\Delta_{c}=5.14\Gamma(5\Gamma)$ for SW(RW) EIT as optimized beam parameters for $\Delta_{p}=5\Gamma$ and center $\omega_{m0}=2\pi\times 3$ MHz. Crosses label points of maximal $W_{c}$. Compared to RW, SW EIT cooling achieves higher cooling rates and lower limits for a larger bandwidth of motional frequency.
Fig. 4 shows the simulated $W_{c}$ and $\bar{n}_{ss}$ as a function of the cooling beam’s Rabi frequency $\Omega_{c}$ and different motional frequencies $\omega_{m}$. The calculations assume a RW pump detuning $\Delta_{p}=5\Gamma$ and repumper amplitude $\Omega_{r}=2\pi\times 10$ MHz in both cases, with $\Delta_{c}$ and $\Omega_{p}$ taken as the optimal values for both RW and SW configurations to target optimal cooling of a motional mode frequency $\omega_{m0}=2\pi\times 3$ MHz. We find an overall improvement using carrier-free EIT in both cooling rate and final phonon number over a significantly expanded parameter range, cooling $\omega_{m}$ between $52\pi\times 1.5-$ 5 MHz to $\bar{n}_{ss}<10^{-1}$.
Figure 5: (a) Time evolution of SW and RW EIT starting at $\langle n\rangle=4$ for $\omega_{m0}=2\pi\times 3$ MHz. Starting at $\bar{n}=4$, SW EIT cooling reaches $\bar{n}_{ss}=0.003$ in $34\,\mu s$ using optimal Rabi frequency $\Omega_{sw}=2\pi\times 42$ MHz for fast cooling rate, while RW EIT takes $75\,\mu s$ to achieve $\bar{n}_{ss}=0.017$ with optimal $\Omega_{rw}=2\pi\times 14$. (b) and (c) Cooling rate and limit for parameters chosen to optimally cool $\omega_{m0}=2\pi\times 3$ MHz as a function of the the pump (b) and cooling (c) beam’s polarization impurity, assumed evenly distributed in the two undesired polarization components in each case, showing the relative insensitivity in cooling beam impurity as compared to the pump. (d) $W_{c}$ and $\bar{n}_{ss}$ for a configuration based on first-order HG delivery of the pump beam (Fig. 3d), as a function of the cooling beam’s TM10 mode power impurity (modeled as fraction of power in the fundamental Gaussian instead of the HG10 mode).
In Fig. 5a, we simulate a GS cooling instance with the optimal Rabi frequencies that correspond to the highest cooling rate at $\omega_{m0}=2\pi\times 3$ MHz in Fig. 4. Starting near the Dopper cooling limit $\bar{n}=4$, carrier-free EIT cools $\bar{n}_{ss}$ to $0.003$, $7$ times lower than that of the RW, within half of the time. Furthermore, operation at lower $\Omega_{c}$ than maximizes $W_{c}$ would allow substantially lower final $\bar{n}_{ss}$ (Fig. 4a) still. These enhancements in the cooling performance, especially in $\bar{n}_{ss}$ over a broader motional mode bandwidth than obtained by RW EIT could considerably alleviate the cooling time overhead by reducing or eliminating the need for further SBC.
For integrated implementations enabling practical carrier-free EIT cooling, delivery of the highly pure circular polarization required to realize the effective three-level system presents a challenge Massai et al. (2022). To guide design of experimental configurations realizing these functions, we quantify polarization impurities’ effects on the predicted performance. We note that implementations at larger magnetic field with larger frequency selectivity can significantly alleviate the required polarization purity Marinelli (2020); however, below we focus on the more challenging polarization purities required for low-field (few G) regimes often used, assuming $B=10$ G in the simulations.
As detailed below, optical access from integrated photonics and $B$-field arrangements impose constraints on design that generally allow practical delivery of one beam with near-perfect polarization purity but compromises in the other’s. We consider the effects of both pump and cooling beams’ polarization impurities separately. Fig. 5b shows cooling rates and limits as a function of the pump beam’s power impurity $\epsilon$, evenly distributed between the other two polarization directions such that the polarization vector can be written as $\sqrt{1-\epsilon}\hat{\sigma}_{+}+\sqrt{\epsilon/2}(\hat{\sigma}_{-}+\hat{\pi})$. Fig. 5c shows the same for the cooling beam’s power impurity. Cooling is clearly strongly sensitive to the strong $\hat{\sigma}_{+}$ pump beam’s impurity: for impurities larger than approximately 1% in relative intensity, the advantage of SW EIT cooling is significantly compromised. Relatively more tolerance holds for the weaker $\hat{\pi}$ cooling beam’s impurity.
These polarization and mode purity requirements inform requirements for potential experimental configurations for implementation of this cooling scheme. While linear polarization purity can straightforwardly be achieved well below 0.1% with integrated photonic addressing Mehta and Ram (2017), $<1\%$ purities are challenging to achieve for circularly polarized emission and present a challenge for design He and Li (2014); Spektor et al. (2023). Fig. 3d gives two possible experimental schemes, both utilizing integrated delivery of the cooling beam but based either on free-space or integrated delivery of the $\hat{\sigma}_{+}$ beam. In the latter case, a vertical B-field is chosen given the likely advantage of a surface-normal-propagating $\hat{\sigma}_{+}$ beam with high purity from passively robust structures Massai et al. (2022); Natarajan et al. (2024); $\pi$-polarized cooling light then requires vertical B-field, which is approximately achieved via grating emission of a TM-polarized waveguide mode at a shallow emission angle propagating close to parallel with the trap chip. The ${\sim}10\%$ tolerance on the $\hat{\pi}$-beam polarization purity allows for emission angles for quasi-TM-polarized cooling beams of approximately $72.5^{\circ}$ from normal.
Naturally, the advantage of carrier-free EIT cooling relies on the stable driving of the ions at the nodal points of the cooling beam. For the three-level system, ref. Zhang et al. (2012) shows that $\bar{n}_{ss}$ degrades by 3×3$\times$3 × for a positional deviation from the cooling beam’s node of $0.017\lambda$ with $\lambda$ the cooling beam’s wavelength. This corresponds to approximately $1\%$ of the SW’s peak anti-node intensity. Ion addressing in a phase-stable SW with a similar level of position accuracy has been demonstrated with integrated photonicsVasquez et al. (2023), indicating that the required positioning precision is practically achievable in present traps with integrated photonics. The bottom design in Fig. 3d utilizes fully integrated beams Massai et al. (2022), with cooling field delivered through TM10 mode. Fig. 5d shows that simulated $\bar{n}_{ss}$ in the eight-level system degrades by an amount consistent with above for 1% power impurity in the fundamental mode, agreeing with Zhang et al. (2012). This sensitivity sets a design goal for experiments to suppress the fundamental mode’s relative intensity to approximately 1% level to avoid unwanted scattering.
Our results indicate significant possible gains in GS cooling via carrier-free EIT cooling, simultaneously in motional mode frequency bandwidth, final phonon number, and cooling rate. By potentially significantly reducing requirements for subsequent SBC, these techniques may alleviate a key bottleneck in trapped-ion system runtime. Polarization purity in pump beam delivery is quantified as a key challenge for implementation at low magnetic fields, along with mode purity requirements when utilizing HG mode drives.
III Conclusion
By selectively driving desired sideband transitions, carrier-free cooling at nodal points of SW or HG modes allows advantages over cooling with running waves in cooling speed and final phonon limit. Modest advantages obtain for Doppler cooling of highly excited ions from well beyond the LDR, with significant enhancements predicted for GS cooling via EIT including in motional mode frequency bandwidth addressed. The proposed schemes utilize either simple SW or HG modes that can be stably and robustly delivered to ions in integrated photonic architectures. Our results indicate the potential for scalable photonics to assist in bypassing key physical limitations of trapped-ion systems, and inform experimental realizations in progress.
IV Acknowledgments
We thank Erich Mueller for helpful discussions. We acknowledge support from an NSF CAREER award (2338897), IARPA via ARO through Grant Number W911NF-23-S-0004 under the ELQ program, the Alfred P. Sloan Foundation, and Cornell University.
Appendix A Lindblad master equation beyond the LDR
The dynamics of a trapped ion in a harmonic potential, interacting with a laser field, can be described using the Lindblad master equation. Inside LDR, Eq. (1) is approximated to first order $\langle e|V_{dip}|g\rangle\approx\frac{\Omega_{0}}{2}(1+i\eta(\hat{a}+\hat{a}^{\dagger}))$, resulting in a carrier and two first-order sideband transitions. The strongest present coupling, carrier transition, governs the internal states of the ion. This allows us to approximate the system by decoupling the external motional states from the internal dynamics and describe the cooling behavior with the rate equation Cirac et al. (1992); Stenholm (1986). These approximations fail when going outside the LDR or when sideband transitions become the leading order coupling as in the carrier-free cooling configurations considered here. Thus, to investigate the dynamics outside the LDR and quantify the enhancement limit of carrier-free cooling, we directly simulate the Lindblad master equation with higher-order terms.
Figure 6: Comparison of simulation (points) and analytical results (lines) for cooling rate over $\omega_{m}$ for the low-intensity limit of cooling beam. The pump beam’s intensity and cooling beam’s detuning are optimized for $\omega_{m0}=2\pi\times 3$ MHz and $\Delta_{p}=5\Gamma$. Blue and red represent results for cooling beam delivered as a SW and RW, respectively. (a) uses a SW pump beam with the ion at an antinode, with $\eta_{c}=\eta_{0}$ and $\Omega_{c}=2\pi\times 2$ MHz for both SW and RW cooling beam. (b) uses RW pump beam. Limited by the selection rule, we choose $\eta_{c}=-\eta_{p}=\eta_{0}/\sqrt{2}$ for RW EIT to maximized $(\eta_{c}-\eta_{p})^{2}$ as in the usual experimental design Lechner et al. (2016). For carrier-free EIT, we choose $\eta_{p}=\eta_{0}/\sqrt{2}$ for the RW pump and $\eta_{c}=\eta_{0}$ for the SW cooling beam for simplicity as only the coupling term $\eta_{c}\Omega_{c}$ matters at SW node in the LDR. The same LD parameters are used in the simulations shown in previous sections. Lastly, we choose $\Omega_{c}=2\pi\times 2\sqrt{2}$ MHz for SW cooling beam and $\Omega_{c}=2\pi\times 2$ MHz for RW cooling beam so that their cooling rates at $\omega_{m0}$ agree in QRT calculation. The presence of pump-beam sideband couplings present in (b) distorts the good agreement between the analytical and master-equation treatment in (a); both treatments show the enhancement in motional frequency bandwidth associated with SW cooling beam delivery.
Higher-order terms in the dipole interaction Hamiltonian come immediately from Taylor expansions as in Eq. (1); higher-order dissipative terms can be derived similarly with some subtleties. For a two-level system with isotropic spontaneous decay rate $\Gamma$, the Liouvillian
[…]
$=$ $\frac{1}{2}\sum_{m}[2\hat{c}_{m}\rho\hat{c}_{m}^{\dagger}-\hat{c}_{m}\hat{c}_{m}^{\dagger}\rho-\rho\hat{c}_{m}\hat{c}_{m}^{\dagger}]$
with $\hat{A}=e^{ixk_{0}\hat{R}}\hat{\sigma}^{-},\hat{\sigma}^{-}=|g\rangle\langle e|$, $x=\cos(\theta)$ for emission angle $\theta$ with respect to motional axis Cirac et al. (1992). Using the Baker–Campbell–Hausdorff formula, $e^{ixk_{0}\hat{R}}=e^{ix\eta_{0}(\hat{a}^{{\dagger}}+\hat{a})}=e^{-\eta_{0}^{2}x^{2}/2}e^{ix\eta_{0}\hat{a}^{{\dagger}}}e^{ix\eta_{0}\hat{a}}$. Denote $\hat{b}_{k}=\frac{(i\eta_{0}\hat{a}^{{\dagger}})^{k}}{k!},\hat{d}_{k}=\frac{(i\eta_{0}\hat{a})^{k}}{k!}$ ($\hat{b}_{j}=\hat{d}_{j}=0\text{ if }j<0\text{ or }j>n$), and Taylor expanding $\hat{A}$ to the $n$th term, we can then calculate the jump operator for the Liouvillian with ±mplus-or-$\pm m$ ($0\leq m\leq n$) change in motional states as
[…]
Appendix B Analytical EIT treatment
Our analytical description of carrier-free EIT cooling follows the approach of Zhang et al. (2012) for an ideal three-level system. A three-level system withpump beam of carrier strength $\Omega_{p0}$ and sideband strength $\eta_{p}\Omega_{p1}$ (coupling states $|r\rangle$ and $|e\rangle$) and cooling beam of carrier strength $\Omega_{c0}$ and sideband strength $\eta_{c}\Omega_{c1}$ (coupling states $|g\rangle$ and $|e\rangle$) can be represented by a rotating-frame Hamiltonian:
$\hat{H}_{0}/\hbar=$ $\nu a^{{\dagger}}a-\Delta_{c}|e\rangle\langle e|-(\Delta_{c}-\Delta_{p})|r\rangle\langle r|$
$+\frac{1}{2}\big{(}\Omega_{c0}\hat{\sigma}_{gx}+\eta_{c}\Omega_{c1}\hat{\sigma}_{gy}\hat{R}+\Omega_{p0}\hat{\sigma}_{rx}+\eta_{p}\Omega_{p1}\hat{\sigma}_{ry}\hat{R}\big{)}$
$\equiv$ $\nu a^{{\dagger}}a+\hat{H}_{\mathrm{int}}-\hat{F}\hat{R}$ (4)
with $\nu$ the motional frequency. We define the force operator $\hat{F}\equiv-\frac{\Omega_{c1}}{2}\eta_{c}\hat{\sigma}_{gy}-\frac{\Omega_{p1}}{2}\eta_{p}\hat{\sigma}_{ry}$, where $\hat{\sigma}_{gy}\equiv-i(|e\rangle\langle g|-|g\rangle\langle e|)$ and $\hat{\sigma}_{ry}\equiv-i(|e\rangle\langle r|-|r\rangle\langle e|)$. The calculation proceeds by approximating the internal state density matrix as the steady-state solution of the optical bloch equations under the action of the internal-state Hamiltonian $\hat{H}_{\mathrm{int}}$, and then considering the effect of the relatively weak forcing terms as a perturbation.
The fluctuation spectrum corresponding to the forcing term is
[…]
$=$ $\lim_{t\rightarrow\infty}\{(\frac{\Omega_{c1}}{2}\eta_{c})^{2}\mathcal{F}[\langle\hat{\sigma}_{gy}(\tau+t)\hat{\sigma}_{gy}(t)]\rangle$ (5)
$+(\frac{\Omega_{p1}}{2}\eta_{p})^{2}\mathcal{F}[\langle\hat{\sigma}_{ry}(\tau+t)\hat{\sigma}_{ry}(t)]$
$+\eta_{c}\eta_{p}\frac{\Omega_{c1}\Omega_{p1}}{4}\big{(}\mathcal{F}[\langle\hat{\sigma}_{gy}(\tau+t)\hat{\sigma}_{ry}(t)\rangle]$
$+\mathcal{F}[\langle\hat{\sigma}_{ry}(\tau+t)\hat{\sigma}_{gy}(t)\rangle]\big{)}\},$
where $\mathcal{F}$ denotes the Fourier transform with respect to $\tau$. Using the quantum regression theorem (QRT) Gardiner and Zoller (2004), we can solve for an analytical expression of (5) from the steady-state solution to $\hat{H}_{\text{int}}$.
When the ion sits at the nodal point, we have $\Omega_{c0}=0$ in the $\hat{H}_{\text{int}}$. Thus, only the first term in (5) is nonzero. Taking the leading order of $\Omega_{c1}$,
$\text{Re}[\mathcal{S}(\nu)]=\frac{2\eta_{c}^{2}\Omega_{c1}^{2}\Gamma(\Delta+\nu)^{2}}{4\Gamma^{2}(\Delta+\nu)^{2}+\big{(}\Omega_{p}^{2}-4(\Delta+\nu)(\Delta_{c}+\nu)\big{)}^{2}},$
where $\Gamma=\Gamma_{c}+\Gamma_{p}$ and $\Delta=\Delta_{c}-\Delta_{p}$. When $\Delta=\nu$, this gives the cooling rate in Zhang et al. (2012) for ideal carrier-free EIT, which does not depend on $\eta_{p}$.
For the standard RW EIT cooling, taking the leading order of $\Omega_{c}$ again, (5) reduces to a simple expression when $\Delta=\Delta_{c}-\Delta_{p}$, proportional to $(\eta_{c}-\eta_{p})^{2}\Omega_{c1}^{2}$. This motivates the counterpropagating cooling and pump beam in experiments Lechner et al. (2016).
The analytical treatment above assumes that sideband couplings do not affect the steady state solution to the optical Bloch equation, valid in the LDR and in the low intensity limit for both $\Omega_{c1}$ and $\Omega_{p1}$. In reality, the effect of the pump beam’s sideband coupling on the internal dynamics is always nonnegligible. Fig. 6 shows cooling rates for for EIT using SW and RW pump beam, comparing the simulation with the analytical results from (5) under the low-intensity limit of cooling beam. For the SW pump beam, we place the ion at its anti-node, where only the carrier is present. This discrepancy between the simulation and analytical results for RW pump beam showcases the necessity of full master simulation in understanding the physical cooling performance.